OpenStudy (nubeer):

find the area of the surface generated by revolving r^2=cos2x about y axis. this question is related to polar coordinates.

OpenStudy (turingtest):

\[r^2=\cos(2x) \]? strange to have r in terms of x like that. If it is in polar coordinates it should be \[r^2=\cos(2\theta)\]though I'm not sure how to revolve that about the y-axis... Please be sure the questions is posted properly.